OpenStudy (anonymous):

Algebra. Solve by elimination method. 2x+3y=6 3x+5y=15

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (watchmath):

Multiply the first equation by -3 and the second by 2 to get -6x-9y=-18 6x+10y=30 Add them to get y=22 Substitute to the first equatin 2x+66=6 2x=-60 x=-30 Therefore (-30,22) is the solution.

OpenStudy (anonymous):

Isn't -18 + 30= 12?

OpenStudy (gw2011):

3x+5y=15 2x+3y=6 Multiply the first equation by 2 and the second by 3 to find y and you get: 6x+10y=30 6x+9y=18 Subract this equation from the first one and you get: y=12 Substitute this value into one of the equations and you get: 2x+3y=6 2x+3(12)=6 2x+36=6 2x=-30 x=-15 Therefore, x=-15 and y=12